an investment broker invests $96,000 in bonds andearns 4% per year on the investment. how much money is earned per year?
Answer by checkley79(3341)

$=P(1+I}^T
$=96000(1+.04)^1
$=96000*1.04
$=99840 ANS. FOR THE 1 YEAR INVESTMENT.